OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2).

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y1-y2}{x1-x2}\) now,just put the values and find m.

OpenStudy (anonymous):

no comprende

OpenStudy (anonymous):

m = slope. y2 and y 1 , x2 and x 1 are all coordinates on your graph

OpenStudy (anonymous):

so in the coordinate (4,8) which is x?

OpenStudy (anonymous):

m= y2-y1 / x2- x1 m=(-2-8) / (6-4) m=-10 / 2 m=-5

OpenStudy (anonymous):

@Kelly226, that does nothing for he user to help him to figure out how to solve the problem. It is much more constructive if you do not just give the answers with work shown or not work shown.

OpenStudy (anonymous):

i showed the steps too.. i didn't just give the answer

OpenStudy (anonymous):

was on cig break whoa calm dwn there horsies lol